OpenStudy (anonymous):

***PLEASE! Help Needed!!?!*** ***Medal AND Fan Rewarded*** - A block of wood weighing 18 pounds rests on an inclined plane that makes an angle of 60 with the horizontal. How much of the friction force is needed to prevent the block from sliding on the plane? - Round answer to nearest whole number.

OpenStudy (ash2326):

Let's make the free body diagram of the block

OpenStudy (ash2326):

 Do you understand the figure and the forces which I have marked

OpenStudy (anonymous):

yes...

OpenStudy (ash2326):

Good, now we need to have friction force such that it cancels the mg sin 60 which is accelerating the block downwards. Let the friction force be F, so to prevent sliding \[F=mg \sin 60\] m= 60 pounds, convert in Kgs g=9/8 or 10 m/s^2 \[\sin 60=\frac {\sqrt 3}{2}\]] Can you solve it now?

OpenStudy (anonymous):

yes... give me a sec to do it...

OpenStudy (ash2326):

ok

OpenStudy (anonymous):

im sorry but i dont comprehend how to do it....

OpenStudy (ash2326):

do you understand that the friction force should be equal to mg sin 60?

OpenStudy (anonymous):

Forget it.... Dont mean to be rude but I dont get this and your not really explaining it in a way that I can comprehend it...

OpenStudy (ash2326):

Ok, let's start again

OpenStudy (ash2326):

Did you understand the free body diagram?

OpenStudy (anonymous):

Can you help me with this problem? Im stuck on the exact one. (A block of wood weighing 18 pounds rests on an inclined plane that makes an angle of 60° with the horizontal. To the nearest pound, how much friction force is needed to prevent the block from sliding on the plane?)